### v2

# Symmetry Breaking and Mie-tronic Supermodes in Nonlocal Metasurfaces

Abstract
Symmetry breaking in metasurfaces is typically used as a means of weakening optical confinement, e.g. transforming infinite $Q$ bound states in the continuum into high $Q$ quasibound states. Within the Mie-tronics framework, we show that symmetry breaking can instead enhance light trapping by strengthening in-plane nonlocal coupling pathways. Through diffraction and multiple-scattering analyses, we demonstrate that diffractive bands and Mie-tronic supermodes originate from the same underlying Mie resonances but differ fundamentally in physical nature. Finite arrays exhibit $Q$-factor enhancement driven by redistributed radiation channels, reversing the trend predicted by infinite-lattice theory. We further show that controlled symmetry breaking opens new electromagnetic coupling channels, enabling polarization conversion in nonlocal metasurfaces. These findings establish a unified wave picture linking scattering and diffraction theories and outline design principles for multifunctional metasurfaces that exploit nonlocality for advanced light manipulation, computation, and emission control.
preprint: APS/123-QED
I Introduction
Optical metasurfaces are rapidly evolving from fundamental science to practical technology, enabling compact platforms for imaging, sensing, energy harvesting, and information processing [Brongersma2025second]. Conventional metasurfaces comprise weakly interacting subwavelength elements, allowing a local description. Yet, recent analyses of their limits show that surpassing these constraints demands nonlocal design [Cai2020inverse, Shastri2023nonlocal, Yao2024nonlocal]. When inter-element coupling becomes strong, collective optical excitations emerge, giving rise to nonlocal metasurfaces governed by coherent, system-scale modes. Because these modes are typically excited by plane waves, they can often be interpreted through diffraction theory, earning the name diffractive nonlocal metasurfaces [Overvig2022diffractive]. Understanding such systems benefits from revisiting the classical diffraction theories that first distinguished local from nonlocal behavior.
In his 1907 “Dynamical Theory of Gratings”, Rayleigh extended Fresnel’s scalar treatment of reflection and diffraction to periodic metallic surfaces. He explicitly noted that Fresnel’s local approximation fails when the grating period approaches the wavelength, observing that the recesses “act as resonators” [Rayleigh1907dynamical]. Rayleigh thus recognized, decades ahead of modern electromagnetic modeling, that resonance fundamentally alters diffraction and that a “more strictly dynamical theory” was required. Fano later provided this framework [Fano1941theory], introducing coupling among diffracted and evanescent orders to describe how these “Rayleigh cavities” form quasi-stationary surface waves that interfere with the continuum, giving rise to asymmetric Fano resonances. The Rayleigh–Fano formalism laid the foundation for modern numerical solvers such as the rigorous coupled-wave analysis (RCWA) or Fourier modal method [Liu2012s4], which we use here to calculate nonlocal diffraction bands in ideal, infinitely periodic metasurfaces.
The design of nonlocal metasurfaces often begins with idealized infinite arrays, where symmetry breaking within the unit cell serves as a versatile tool for tailoring the dispersion and radiation properties of diffractive bands [Nguyen2018symmetry, Overvig2020selection, You2024resonance, Barreda2025bound]. Engineering such diffractive bands has recently fueled advances in analog optical computing and wave-based information processing [Guo2018photonic, Kwon2018nonlocal, Komar2021edge, Liu2024edge]. In practical implementations, however, metasurfaces are finite—a feature that has drawn growing attention [Ho2024finite, Karavaev2025emergence, Chen2025observation, Zhang2025vortex]. While compactness makes them appealing for integration, it simultaneously invalidates the assumption of infinite periodicity inherent to diffraction theory. As the array size decreases, translational symmetry is broken and the momentum conservation underlying the Fourier modal method no longer holds. The resulting collective resonances become discrete and strongly dependent on the global geometry of the structure. In this finite regime, multiple scattering and modal hybridization provide a more accurate physical description [Lagendijk1996resonant, Hoang2025collective].
At the microscopic level, these collective effects originate from interactions among Mie modes—the localized resonances of individual dielectric or metallic elements—which constitute the building blocks of Mie-tronics [Rybin2024metaphotonics]. Beyond governing the response of isolated particles [Li2024machine, Barati2025toward], the coupling and interference of Mie modes across arrays yield bonding and antibonding supermodes. These supermodes can enhance device performance and enable entirely new metadevice concepts [Krasnok2012all, Babicheva2024mie, Ha2024optoelectronic, Mao2025lateral].
In this work, we systematically investigate how symmetry breaking influences both diffractive and Mie-tronic supermodes. Together with the accompanying Letter [Hoang2025unconventional], we show that, contrary to expectations based on infinite-array diffraction bands, symmetry breaking in finite metasurfaces can enhance in-plane multiple scattering and thereby increase the quality factor ($Q$) of certain nonlocal modes. We further demonstrate that coupling among Mie modes produces additional bonding and antibonding supermodes beyond the Fourier-mode couplings of classical diffraction theory, expanding the functional landscape of nonlocal metasurfaces. Notably, while antibonding supermodes remain robust in the presence of a quartz substrate, bonding supermodes are suppressed. Finally, we show that distinct unit-cell geometries can support identical Mie-resonant supermodes, underscoring the robustness and design versatility of the Mie-tronic framework for both diffractive and finite-size nonlocal metasurfaces.
II Fundamentals of Mie-tronics
II.1 Historical Origins of Mie-tronics
The intellectual roots of Mie-tronics trace back to the 19th-century work of Clebsch, who sought a rigorously wave-based framework for optical devices at a time when lens and mirror design was dominated by geometrical principles [Logan1965survey]. Employing both plane-wave and point-like sources, he introduced an early form of the multipole expansion to solve boundary conditions—an idea that became a cornerstone of scattering theory. Although Clebsch ultimately declared his effort a failure, his pioneering use of multipole expansions endured [Devaney1974multipole]. What once seemed unattainable is now well established: the laws of reflection and refraction follow directly from wave physics [Bohren2008absorption], specifically from the scattering coefficients obtained at a spherical interface [Hoang2012interpretation, Hoang2013rigorous]. In this sense, Clebsch’s attempt to reinterpret bulky optical elements in wave terms resonates with today’s drive to realize ultrathin metasurfaces that harness wave phenomena to replace conventional optics.
Clebsch’s ideas influenced Lorenz, who first derived the exact solution for plane-wave scattering by a sphere. This result, however, later became more widely associated with Mie, whose systematic multipole analysis laid the analytical foundation of modern scattering theory [Wriedt2012mie]. Unlike Lorenz, Mie was primarily inspired by Rayleigh, who had independently investigated acoustic and optical scattering problems similar to those considered by Clebsch. Notably, Mie introduced the now-standard term “Rayleigh scattering” to distinguish the point-particle regime from his treatment of larger spheres [Mie1976contributions].
Over the following century, these contributions coalesced into the framework of generalized Lorenz–Mie theory, emphasizing the dual importance of intrinsic scatterer modes and the multipole content of excitation sources [Gouesbet2011generalized]. Advances in nanoscience have since extended scattering analysis to particles of arbitrary shape [Kildishev2025art], enabling entirely new classes of photonic devices [Dorodnyy2023mie]. This convergence of multipole physics and nanophotonics defines the emerging paradigm of Mie-tronics [Won2019into]—a framework that unifies classical scattering theory with modern metasurface design.
II.2 Diffraction and Scattering Frameworks
The early development of scattering theory was closely intertwined with the study of diffraction. In the 19th century, Clebsch, Lorenz, and Rayleigh often described light–particle interactions in the language of diffraction [strutt1871lviii]. Before Maxwell’s electromagnetic formulation, when the Huygens–Fresnel principle dominated, the terms “diffraction” and “scattering” were used almost interchangeably. With the establishment of Maxwell’s equations, however, diffraction came to be recognized as a specific manifestation of the broader phenomenon of electromagnetic scattering [Rayleigh1881x, Feynman2015feynman].
Even today, these terms persist in distinct theoretical contexts, sometimes leading to ambiguity. For example, in diffractive nonlocal metasurfaces, “scattering” is often used to describe the radiative leakage of bound states that fundamentally originate from diffraction processes [Hoang2025collective]. Classical diffraction theory typically considers coupling through a single Fourier order as the primary radiation channel, but electromagnetic waves naturally support multiple scattering pathways due to their positive-definite energy spectrum. Recognizing these additional channels, beyond the standard Fourier harmonics, is essential for accurately describing nonlocal diffraction bands [Feng2025beyond] and for engineering superlattices with tailored spectral responses [Song2025emergence]. In formal scattering theory, enhancing the $Q$ of nonlocal modes corresponds to strengthening constructive multiple scattering rather than suppressing it.
For conceptual clarity, we adopt the following convention: Results obtained using RCWA—where in-plane periodic boundary conditions are imposed and electromagnetic fields are expanded into Fourier harmonics under plane-wave excitation—are referred to as diffraction theory, rooted in the classical grating analyses of Rayleigh and Fano [Rayleigh1907dynamical, Fano1941theory]. In contrast, results derived from multipole expansions of scattered fields in finite metasurfaces, solved under open (radiative) boundary conditions, are referred to as scattering theory, grounded in the pioneering works of Clebsch, Lorenz, Rayleigh, and Mie [Logan1965survey].
As shown below, these two perspectives—diffraction in infinite periodic systems and scattering in finite structures—merge naturally within the Mie-mode basis. Within this unified framework, Mie-tronics bridges Fourier-mode analyses with multipole expansions, revealing their common physical foundations and providing a coherent language for describing nonlocal metasurfaces.
II.3 Mie Coefficients as Building Blocks
Figure 1: Beyond spheres: unit cells in Mie-tronics. (a) Schematic of a magnetic dipole ($m_{x}$) interacting with an array of air holes in a silicon slab. Insets show two unit cells that are equivalent from a photonic-crystal perspective but distinct in Mie-tronics, with the geometrical parameters indicated. (b) Four functionally equivalent unit cells and their parameters. The spherical unit cell has a period of 425 nm, while the square and T-shaped unit cells have a period of 720 nm. All unit cells are composed of silicon ($n=3.5$) and surrounded by vacuum.
Although practical scattering systems are more complex than a single sphere, Mie coefficients and multipole expansions remain fundamental tools for understanding their optical behavior. We therefore briefly review the standard Mie formalism, its implementation, and physical meaning. In this framework, the incident ${\bf E}^{\text{inc}}$, scattered ${\bf E}^{\text{sct}}$, and internal ${\bf E}^{\text{int}}$ fields are expanded in a multipole basis:
${\bf E}^{\text{inc}}({\bf r})$ $=\sum_{l=1}^{L_{\text{i}nc}}\sum_{m=-l}^{l}\left[p_{lm}{\bf N}_{lm}(k{\bf r})+q_{lm}{\bf M}_{lm}(k{\bf r})\right],$
${\bf E}^{\text{sct}}({\bf r})$ $=\sum_{l=1}^{L_{\text{R}}}\sum_{m=-l}^{l}\left[a_{l}p_{lm}{\bf N}^{(1)}_{lm}(k{\bf r})+b_{l}q_{lm}{\bf M}^{(1)}_{lm}(k{\bf r})\right],$
${\bf E}^{\text{int}}({\bf r})$ $=\sum_{l=1}^{L_{\text{R}}}\sum_{m=-l}^{l}\left[c_{l}p_{lm}{\bf N}_{lm}(nk{\bf r})+d_{l}q_{lm}{\bf M}_{lm}(nk{\bf r})\right],$
where $a_{l}$ and $b_{l}$ ($c_{l}$ and $d_{l}$) are the external (internal) Mie coefficients, determining how each multipole component radiates into free space or induces displacement currents within the sphere. Here $n$ is the sphere’s refractive index and $k=2\pi/\lambda$ the free-space wavenumber.
When the incident field has an analytical multipole representation with order $L_{\text{inc}}$, the solution is fully analytical. In the original Mie formulation, however, the series extends to infinity and must be truncated at a finite order $L_{R}$, yielding a semi-analytical solution. An empirical rule for a sphere of radius $R$ is [Hoang2017fano]:
$L_{\text{R}}\approx 2\pi R/\lambda+4.05\sqrt[3]{2\pi R/\lambda}+2.$
For example, the sphere in Fig. 1 requires $L_{\text{R}}$ between 7 and 10 for wavelengths near 1500 nm. While our focus is on magnetic dipole (MD, $l=1$) interactions, setting $L_{\text{R}}=1$ is inadequate because it neglects higher-order coupling, leading to inaccurate estimates of $Q$ factors and Purcell enhancements. Moreover, when $L_{\text{i}nc}>L_{\text{R}}$, the higher-order incident components ($l>L_{\text{R}}$) cannot couple to internal Mie modes and thus propagate unscattered.
Mie coefficients are formally valid only for isotropic, homogeneous spheres. For arbitrary geometries—such as the hole-type, square, and T-shaped unit cells in Fig. 1—full-wave simulations are required. Here we employ the finite-difference time-domain method, which agrees closely with analytical Mie theory [Hoang2025collective]. For general arrays, Mie-tronics provides a physically meaningful multipole basis for analyzing and interpreting collective resonances.
Figure 1(a) shows a typical photonic-crystal slab interacting with an out-of-plane magnetic dipolar emitter, producing a vortex beam [Chen2025observation], identified as a bound state in the continuum (BIC). From a Mie-tronics perspective, the unit cell in the right inset of Fig. 1(a) is functionally equivalent to the three other unit cells in Fig. 1(b). Arrays of these geometries—adapted from previous works [Chen2022observation, Liu2019high, Hoang2024photonic]—support the same collective Mie resonances despite their differing shapes.
Before addressing other geometries, the remainder of this section examines the spectral profiles of Mie coefficients and how supermodes (collective resonances) arise in finite arrays of spheres, as well as their connection to Bloch-band formation.
Figure 2: Spectral dependence of the lowest-order Mie coefficients of the silicon sphere in Fig. 1. Internal (left) and external (right) electric-dipole (ED) and magnetic-dipole (MD) coefficients reveal broad overlapping resonances. Values exceeding unity in the internal coefficients indicate enhanced energy storage within the particle, consistent with whispering-gallery–type behavior.
Figure 2 shows the spectral profiles of the two lowest-order Mie coefficients: the electric (ED) and magnetic dipoles. The internal ED and MD resonances exhibit broad spectral distributions. For a sphere with unit index contrast, the internal coefficients equal 1; thus, values exceeding unity indicate enhanced energy storage. The ED mode spans nearly the entire 1200 nm range, while the MD mode extends from 1200 to 1940 nm. Such broad features enable hybridization among distinct Mie modes, giving rise to diverse collective resonances. They also explain why analytical models restricted to a single Mie mode fail to capture realistic $Q$ factors and Purcell enhancements [Hoang2024photonic].
Despite their broad spectra, the ED and MD resonances correspond to whispering-gallery modes, consistent with both geometrical and wave-optics descriptions. Early photonic-crystal studies argued that large unit cells would not support photonic bands due to the dominance of geometrical optics [John1991localization]. Mie-tronics revises this view: all Mie modes inherently embody wave physics and can couple collectively. As shown in prior work, coupling between low- and high-order modes gives rise to photonic band formation [Hoang2017fano, Hoang2020collective].
Spectral overlap among Mie modes thus lays the foundation for collective optical phenomena in arrays. When resonant unit cells are periodically coupled, their individual multipolar responses hybridize to form delocalized collective modes—Mie-tronic supermodes—in finite arrays. In the following section, we examine how these collective Mie-tronic supermodes connect to Bloch bands in infinite periodic systems.
II.4 Beyond Bloch Bands: Antibonding and Bonding Mie-tronic Supermodes
Figure 3: Mie-tronic origin of supermodes in finite arrays and their connection to Bloch bands in photonic crystals. (a) Purcell-factor spectrum for the configuration in the inset, revealing two well-separated bands (antibonding and bonding) with pronounced supermodes labeled $A_{1,2}$ and $B_{1}$. (b),(c) Magnetic ($H_{x}$) and electric ($E_{z}$) field distributions for the antibonding mode $A_{1}$; the inset shows the corresponding far-field radiation pattern. (d) Bloch bands associated with the antibonding and bonding supermodes in (a). (e),(f) Same as (b),(c), but for the bonding mode $B_{1}$; the inset highlights dominant in-plane radiation leakage.
Figure 3(a) presents the Purcell-factor spectrum for the configuration shown in the inset, where a magnetic dipole $m_{x}$ couples to a $5\times 5$ array of spheres. Two distinct bands appear, each containing well-defined supermodes labeled $A_{1}$, $A_{2}$, and $B_{1}$. The $A_{1}$ and $B_{1}$ modes correspond to the $\Gamma$- and $M$-point edges of the antibonding ($A$) and bonding ($B$) Bloch bands in Fig. 3(d), which emerge in the limit of an infinite array. To obtain the dispersions in Fig. 3(d), the dipole $m_{x}$ is positioned at the sphere center, and Bloch boundary conditions are applied along the $y$ and $z$ directions.
Because the multipole modes are orthogonal, the source initially excites only the intrinsic MD resonance, for which the internal field at the sphere center is determined by the MD amplitude $\eta_{x}=\sqrt{\tfrac{3}{8\pi}}\lambda H_{x}$, where $H_{x}$ is the local magnetic field [Hoang2022high]. Optical coupling across the Bloch boundaries subsequently induces higher-order multipoles, capturing multiple-scattering processes in the lattice. As a result, the calculated band-edge wavelengths align precisely with the observed supermodes [Hoang2020collective].
The antibonding and bonding characters of these modes are evident from their near- and far-field profiles. Figures 3(b,c) show the $A_{1}$ (antibonding) supermode, while Figs. 3(e,f) show the $B_{1}$ (bonding) mode. The in-plane electric-field ($E_{z}$) distributions reveal that $A_{1}$ exhibits nodes between adjacent spheres, whereas $B_{1}$ displays pronounced hotspots within the interparticle gaps. These hotspots signify stronger in-plane coupling and larger Purcell enhancement for the bonding mode. Such enhanced local fields enable efficient interaction with emitters or detectors oriented along the $z$ direction, suggesting on-chip architectures where Mie-tronic modes act as optical communication channels linking integrated transmitters and receivers—forming scalable photonic networks capable of emulating quantum many-body interactions. The topological properties of Mie modes, in this context, could underpin future classical and nonclassical simulators and light sources [Nielsen2010quantum, Berloff2017realizing, Sergeeva2025coherent, Deng2025enhancement].
A further distinction arises from the orientation of the magnetic moments. In $A_{1}$, the moments are aligned, leading to repulsive interactions; in $B_{1}$, they are anti-aligned, resulting in attractive coupling. Put differently, bonding supermodes extend their fields into the environment, whereas antibonding modes confine them within the unit cells. This difference explains the greater robustness of antibonding modes to perturbations introduced by a quartz substrate, as discussed in the next section.
The far-field radiation patterns [insets of Figs. 3(b,e)] provide complementary evidence. The bonding mode $B_{1}$ radiates mainly in-plane, while the antibonding mode $A_{1}$ emits strongly out of plane. Similar to the hole arrays discussed in the accompanying Letter [Hoang2025unconventional], the $A_{1}$ mode produces an out-of-plane vortex beam, confirming its collective MD origin. From an applications perspective, the in-plane confinement of $B_{1}$ makes it especially appealing for on-chip photonic devices that demand minimal out-of-plane losses.
Interestingly, neither $A_{1}$ nor $B_{1}$ radiates along the surface normal ($x$ axis), precluding excitation by normally incident plane waves. This symmetry protection follows directly from their near-field profiles. A normally incident $z$-polarized plane wave can excite only modes with $E_{z}$ even in zz—a condition unmet by either $A_{1}$ or $B_{1}$. Similarly, a $y$-polarized wave excites only modes with $E_{y}$ even in $y$, but here $E_{y}$ is odd, again forbidding coupling. The same constraints can also be viewed as the parity of the Bloch eigenfields within the unit cell. These symmetry considerations offer a useful guideline for tailoring diffractive bands through controlled symmetry breaking, as we demonstrate in the following section.
Nevertheless, the collective nature of Mie-tronic supermodes allows normal-incidence excitation in sufficiently large arrays [Hoang2025collective]. In Fig. 3(a), only three discrete supermodes appear for the $5\times 5$ array, but their number increases rapidly with array size. Retardation introduces additional supermodes whose $E_{z}$ distributions become symmetric about the $z$ axis, enabling coupling to normally incident plane waves. The discrete peaks in Fig. 3(a) arise from the radiative boundary conditions of finite arrays, capturing high-$Q$ supermodes inaccessible to conventional Bloch-wave analysis, which assumes in-plane periodicity and thus yields only continuous bonding and antibonding bands [Fig. 3(d)].
These finite-array effects highlight how the interplay between collective coupling and structural symmetry profoundly reshapes resonance behavior—a theme we now pursue by examining how symmetry breaking governs the formation and quality factors of Mie-tronic supermodes.
III Impact of Symmetry Breaking on Mie-tronic Supermodes
Symmetry breaking in unit cells has been extensively explored for engineering diffractive bands, yet its influence on collective resonances in finite metasurfaces remains far less understood. Here, we show that symmetry breaking not only modifies the coupling between Mie-tronic supermodes and external radiation but can also produce counterintuitive $Q$-factor enhancement—opposite to trends predicted by infinite-lattice Bloch theory. Moreover, the distinct topological characteristics of Mie modes provide a natural foundation for designing nonlocal metasurfaces with tailored confinement, coupling, and polarization properties.
III.1 Robustness of Antibonding Supermodes Against Symmetry Breaking
Figure 4: Symmetry breaking enhances in-plane multiple scattering, preserves antibonding supermodes, and suppresses bonding counterparts in square and T-shaped arrays. (a) Purcell-factor spectrum for a magnetic dipole $m_{x}$ at the center of a $9\times 9$ array of square unit cells. The presence of a quartz substrate ($n\simeq 1.45$) suppresses the bonding mode $B_{1}$, while the antibonding mode $A_{1}$ remains nearly unaffected. (b),(c) Magnetic near-field distribution and vortex-like far-field pattern of $A_{1}$. (d)–(f) Same as (a)–(c) but for the T-shaped array. The far-field pattern highlights enhanced in-plane multiple scattering along in-plane channels (ICs) versus suppressed out-of-plane channels (OCs).
Figure 4 summarizes the effects of in-plane and out-of-plane symmetry breaking on Mie-tronic supermodes. In-plane symmetry breaking is introduced by replacing square unit cells with T-shaped ones (same lattice vectors but reduced unit-cell symmetry), while out-of-plane breaking is implemented by adding a quartz substrate. For the free-standing square lattice, two well-separated resonant bands appear [Fig. 4(a)], corresponding to the antibonding ($A_{1}$) and bonding ($B_{1}$) supermodes at the band edge. The magnetic-field distribution of $A_{1}$ [Fig. 4(b)] confirms its MD nature, consistent with the sphere array in Fig. 3. Introducing a substrate strongly perturbs the bonding mode $B_{1}$, whose attractive field distribution extends into the environment and overlaps with substrate-supported channels. In contrast, the antibonding mode remains largely unaffected owing to its repulsive character, which minimizes near-field interaction with the substrate.
The vortex-like far-field pattern of $A_{1}$ [Fig. 4(c)] indicates that its primary radiation leaks out of plane, with weaker components along $\pm y$ and $\pm z$. Compared with the $5\times 5$ sphere array in Fig. 3, the $9\times 9$ square array supports more radiation channels, reflecting stronger in-plane multiple scattering in larger systems. This enhanced scattering manifests as a higher Purcell factor in Fig. 4(a), signifying greater light–matter coupling efficiency.
Figures 4(d)–(f) present the response of the T-shaped array. Transforming the square unit cell into a T shape preserves the lattice but breaks its in-plane symmetry, producing a blue shift in the Purcell spectrum [Fig. 4(d)]. Both $A_{1}$ and $B_{1}$ modes persist in the free-standing configuration, and their substrate responses mirror those of the square array. The magnetic near field [Fig. 4(e)] and far-field pattern [Fig. 4(f)] highlight a crucial effect of in-plane symmetry breaking: radiation is redistributed, with significant leakage redirected from out-of-plane (OC) to in-plane channels (ICs) along $y$ and $z$. This redistribution increases the dwell time of photons within the metasurface, enhancing $Q$ and setting the stage for the results discussed next in Fig. 5.
III.2 $Q$ Enhancement via Symmetry Breaking
Figure 5: Symmetry breaking lowers the $Q$ of Bloch modes in infinite lattices but enhances the $Q$ of supermodes in finite metasurface arrays. (a) Antibonding Bloch bands above the light line for the four unit-cell types. (b) $Q$ of Bloch modes at $\Gamma$: symmetry breaking reduces the divergent $Q$ of the square unit to a finite value in the T-shaped unit, while sphere and hole lattices show much higher $Q$. (c) $Q$ of supermodes in finite arrays: symmetry breaking boosts the $Q$ of T-shaped arrays beyond that of square arrays for sizes $5\times 5$ to $33\times 33$, with both outperforming sphere and hole arrays.
Figure 5(a) shows the antibonding Bloch bands above the light line for the four unit-cell geometries. The calculated band-edge wavelengths agree well with the corresponding $A_{1}$ supermodes (the hole-array case is detailed in the accompanying Letter [Hoang2025unconventional]). Modifying the unit-cell geometry shifts the band edges and alters the dispersion; notably, the square and T-shaped lattices exhibit narrow spectral widths ($\sim$20 nm), nearly an order of magnitude smaller than that of the hole lattice (179 nm). This result highlights the capability of geometrical design to engineer band dispersion with high precision.
Near the $\Gamma$ point [Fig. 5(b)], the Bloch-mode $Q$ diverges for the sphere, square, and hole lattices, reflecting at-$\Gamma$ BICs. In contrast, the T-shaped lattice exhibits a finite $Q$, consistent with the conventional picture in which unit-cell symmetry breaking introduces radiative leakage and reduces confinement [Koshelev2018asymmetric].
Remarkably, the opposite trend emerges in finite metasurface arrays: symmetry breaking enhances optical confinement through stronger in-plane multiple scattering. As shown in Fig. 5(c), the T-shaped array achieves the highest $Q$ across sizes from $5\times 5$ to $33\times 33$, reversing the behavior predicted by infinite-lattice Bloch analysis. The enhancement arises from radiation redistribution: whereas the square array primarily leaks energy out of plane, the T-shaped array redirects radiation into in-plane channels, increasing dwell time by more than an order of magnitude relative to the hole and sphere arrays. For very large arrays, however, the proliferation of radiation pathways eventually saturates the $Q$ enhancement. This interplay between broken symmetry and finite-size effects underscores the need to move beyond Bloch-wave descriptions when designing nonlocal metasurfaces for high-$Q$ performance.
Bloch waves have long served as a powerful framework for describing electron behavior in crystals, underpinning much of the success of the modern electronics industry. However, it has been recognized since shortly after Bloch’s original formulation that, while Bloch waves offer an intuitive picture of electron transport, they neglect the atomic-scale structure of matter and therefore cannot describe electronic localization. This limitation led to the development of Wannier functions [Wannier1962dynamics], which now form the foundation of electronic-structure calculations for finite systems [Marrazzo2024wannier]. Although optical analogues of Wannier functions have been proposed for modeling light localization, the exponential confinement intrinsic to electronic Wannier functions is not directly applicable to electromagnetic waves. As shown in the accompanying Letter [Hoang2025unconventional], multipole functions provide a more natural and physically rigorous basis for investigating light localization within the Mie-tronic framework.
IV Diffractive Nonlocal Metasurfaces
The ongoing revolution in artificial intelligence has renewed interest in analog optical computing, driven by its potential for massive parallelism and ultralow energy consumption. Optical information processing, however, is far from new—it has a long history in optical encryption, image processing, and early forms of optical computing [Ambs2010optical, Chen2014advances]. Around the turn of this century, digital optical computing drew significant attention, but progress was ultimately limited by the weak optical nonlinearities of conventional materials, which prevented the realization of optical logic gates.
The present resurgence stems from the recognition that artificial intelligence operates through statistical rather than logical principles—tasks inherently suited to analog optical systems [Mcmahon2023physics, Zhang2025all]. Within this context, diffractive nonlocal metasurfaces have emerged as a promising platform for high-dimensional analog computation. Their ability to sustain collective, system-scale optical modes enables both deep miniaturization of on-chip architectures and performance that surpasses conventional optical components [Guo2020squeeze, Miller2023optics, Wan2024multidimensional, Javidi2025roadmap].
IV.1 Mie-Resonant Origin of Nonlocal Diffractive Bands
Figure 6: Mie-resonant origins of diffractive nonlocal bands. (a) Schematic of an S- or P-polarized plane wave incident on a photonic-crystal slab. (b) Reflectance of an S-polarized plane wave on the sphere metasurface, revealing a diffractive band. The inset shows the magnetic field component $|H_{x}|$ at an incident angle of $5^{\circ}$. The diffractive nonlocal band vanishes at normal incidence due to symmetry mismatch between the incident field and the supermode. (c) Same as (b) but for the hole metasurface. The inset highlights that the diffractive resonant band originates from the MD resonance.
Diffractive nonlocal metasurfaces arise from the interaction between an incident plane wave and an idealized infinite lattice, as schematically shown in Fig. 6(a). The plane wave coherently drives all unit cells in phase, exciting their intrinsic multipoles simultaneously. Each unit cell acts as a nanoscale oscillator supporting discrete multipolar resonances—most prominently ED and MD—whose spectral overlap is shown in Fig. 2. Higher-order multipoles correspond to higher natural frequencies.
For an isolated sphere, the Cartesian multipoles along the $x$, $y$, and $z$ directions are spectrally degenerate. Coupling within an array lifts this degeneracy, splitting each vectorial multipole into several collective modes. For conceptual clarity, we focus on the $x$-oriented MD resonance throughout this work, although dipoles oriented along $y$ or $z$ can form distinct nonlocal bands under appropriate excitation.
Multipoles of higher order ($l\geq 3$) produce collective modes whose Bloch bands lie entirely above the light line [Hoang2025collective]. Within diffraction theory, such higher-order bands couple to multiple radiation continua, making them particularly relevant for diffractive nonlocal metasurfaces. Their spectra, however, are complex and densely packed. Consequently, most analyses—including the present one—focus on the two lowest-order multipoles, MD and ED, whose well-separated resonances allow controlled dispersion engineering through unit-cell geometry.
In coupled arrays, both dipole and quadrupole modes split into bonding and antibonding branches. Within diffraction theory, only the antibonding modes couple efficiently to the radiation continua, giving rise to observable diffractive bands. Because dipole and quadrupole resonances occur at relatively low frequencies, their associated diffraction bands are spectrally well separated and tunable through geometry. Figures 6(b) and (c) illustrate this effect: reshaping the spherical unit cell into a hole-type geometry transforms the MD-based metasurface from reflection-dominated to transmission-dominated operation. The near-field maps obtained from RCWA (insets) confirm that these diffractive bands originate from the MD resonance.
A key feature of these diffractive bands is their disappearance at normal incidence ($\theta=0^{\circ}$ in Figs. 6(b),(c)), commonly referred to as symmetry-protected BICs. Although these BICs are typically associated with singly degenerate diffraction bands, their physical origin is often left unexamined. Mie-tronics provides a clear interpretation by analyzing how electromagnetic coupling occurs through either electric or magnetic field components [Awai2007coupling].
For the out-of-plane MD mode, magnetic coupling with a P-polarized wave is forbidden because the magnetic field of P polarization lies entirely in-plane and thus remains orthogonal to the out-of-plane MD moment. Electric coupling is likewise prohibited: the in-plane electric field of the MD mode is circularly symmetric and antisymmetric with respect to the P-polarization axis ($z$), as shown in Fig. 3(c), yielding a vanishing net coupling coefficient. Symmetry breaking can lift this restriction, as shown later for the T-shaped unit cell.
In contrast, an S-polarized wave possesses a nonzero magnetic-field component along the out-of-plane direction ($x$) for all oblique incidence angles. Consequently, S polarization can excite the MD resonance except at normal incidence ($\theta=0^{\circ}$), where the out-of-plane magnetic component vanishes. This symmetry argument explains the polarization selectivity of the MD resonance and clarifies the physical origin of symmetry-protected BICs in diffractive nonlocal metasurfaces.
More fundamentally, the symmetry mismatch arises from destructive interference of multipolar displacement currents within the symmetric unit cell, which cancels the radiation of the MD supermode along the surface normal [Hoang2025collective]. This interference picture also explains how symmetry breaking restores coupling between P-polarized plane waves and the MD supermode, as demonstrated in the following subsection.
IV.2 Symmetry Breaking Enables Polarization Conversion
Figure 7: Symmetry breaking enables polarization conversion in nonlocal metasurfaces. (a),(b) Diffractive nonlocal bands in square and T-shaped metasurfaces under S-polarized illumination. Insets show $|H_{x}|$ at an incident angle of $5^{\circ}$; both bands vanish at normal incidence. (c),(f) Transmittance spectra for S-polarized excitation of the T-shaped metasurface, revealing partial conversion from S to P polarization. (d) Reflectance of a P-polarized wave incident on the square metasurface shows no diffractive nonlocal band. (e) Same as (d) but for the T-shaped metasurface, demonstrating that symmetry breaking allows MD resonances to couple to P-polarized waves even at normal incidence.
Breaking in-plane symmetry introduces new coupling channels that relax polarization selection rules. Figures 7(a)–(b) compare the S-polarized reflectance of square and T-shaped metasurfaces, whose near-field distributions (insets) confirm that the diffraction bands originate from the MD resonance. The square unit cell behaves similarly to the spherical case: around each resonant peak, the S-polarized reflectance ($R_{s}$) exhibits a rapid variation, and the diffraction band vanishes at normal incidence due to symmetry mismatch, as discussed earlier for the sphere and hole arrays.
Interestingly, despite the broken in-plane symmetry, the diffraction band of the T-shaped array [Fig. 7(b)] still disappears at normal incidence. Because the T-shaped lattice cannot sustain collective resonances with infinite $Q$ in either finite or infinite configurations [Fig. 5], the vanishing response cannot originate from a zero-linewidth mode. Instead, no resonance is excited at $\theta=0^{\circ}$, which can be understood from symmetry considerations: the T-shaped unit cell breaks symmetry only along the $z$ axis (the magnetic-field direction of S polarization) while retaining mirror symmetry along the $y$ axis (the electric-field direction). Consequently, the eigenmode’s electric component, antisymmetric along $y$, and its magnetic component, oriented along $x$, both remain incompatible with those of normally incident S-polarized light, preventing coupling.
The disappearance of the diffraction band as the incident angle approaches zero is often attributed to the at-$\Gamma$ BIC shown in Fig. 5(b) for symmetric unit cells. However, similar behavior has also been observed experimentally in finite nonlocal metasurfaces [Jie2025near], suggesting that such attribution may be misleading. Because finite structures cannot support true at-$\Gamma$ resonances with infinite QQ—a condition that strictly requires perfect periodicity—the distinction lies in their physical origin: symmetry mismatch does not require an infinite lattice, whereas the at-$\Gamma$ BIC arises from the singularity of wave summation in a perfectly periodic array [Hoang2025collective].
The finite $Q$ of the T-shaped metasurface, discussed in Fig. 5 for both Bloch and scattering waves, is further illustrated in Figs. 7(d) and (e), which compare the P-polarized reflectance of the square and T-shaped lattices. The square array exhibits no diffractive resonance under P polarization due to the same symmetry constraints discussed earlier. In contrast, the T-shaped array enables electric coupling between the eigenmode and the incident P-polarized wave: the eigenmode’s electric field is no longer antisymmetric along the $z$ axis (the electric-field direction of P polarization). This broken symmetry thus allows excitation of the resonant diffraction band even at normal incidence, as shown in Fig. 7(e).
Another notable feature in Figs. 7(b) and (e) is that, across the resonance, the reflectance does not drop from unity to zero, as in symmetric metasurfaces. This behavior indicates the participation of an additional transmission channel in the diffraction process. Because the lattice period is subwavelength, the only accessible channel is the zeroth-order transmission of P polarization. In other words, the T-shaped metasurface partially converts S-polarized incident light into P-polarized transmission. This polarization conversion, confirmed by the spectra in Figs. 7(c) and (f), reaches up to 30%. The conversion efficiency depends on the degree of symmetry breaking, which governs the electromagnetic coupling between the incident wave and the underlying nonlocal modes. These results show that symmetry considerations within the Mie-tronics framework provide a robust strategy for designing nonlocal metasurfaces with tunable polarization filtering and conversion.
V Discussion and Conclusion
Metasurfaces have become a cornerstone of modern photonics, offering miniaturized, chip-scale platforms for advanced optical functionalities. Yet, pushing their performance beyond current limits requires explicit consideration of nonlocal interactions among unit cells. Within the Mie-tronics framework, we have shown that symmetry breaking enables precise control over the coupling between metasurface supermodes and external excitations, leading to polarization conversion and enhanced light confinement in finite metastructures. These mechanisms lay the foundation for emerging functionalities such as analog optical computing and on-chip information processing.
In this unified picture, each unit cell functions as an electromagnetic oscillator supporting multiple multipolar resonances that interact through both short- and long-range electromagnetic coupling [Rider2022advances, Li2025realization, Arjas2025topological]. Advances in high-performance computing now allow direct numerical solutions of Maxwell’s equations for large ensembles of Mie scatterers [Hoang2025unconventional], bridging microscopic multipole physics with macroscopic diffraction phenomena. As metasurface research increasingly embraces nonlocality, the Mie-mode basis offers a universal language for describing and optimizing light–matter interactions beyond Fourier harmonics [Da2025dynamics]. Extending the Mie-tronics framework to nonlinear, quantum, and time-varying regimes will further expand the frontier of nonlocal photonics, paving the way toward multifunctional integrated photonic circuits that seamlessly unite computation, control, and communication within a single platform.
Acknowledgements.
This research is supported by the National Research Foundation, Singapore, and A*STAR under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0009), the Australian Research Council (Grant No. DP210101292), and International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (Contract FA520923C0023). D.L. acknowledges a support from the Ministry of Education of Singapore under its SUTD Kickstarter Initiative (Grant No. SKI 20210501). J. Ji acknowledges the funding received from the PhotonDelta National Growth Fund programme.

### v3

@@
-Symmetry Breaking and Mie-tronic Supermodes in Nonlocal Metasurfaces
+Mie-tronics supermodes and symmetry breaking in nonlocal metasurfaces
 Abstract
-Symmetry breaking in metasurfaces is typically used as a means of weakening optical confinement, e.g.
-transforming infinite $Q$ bound states in the continuum into high $Q$ quasibound states.
-Within the Mie-tronics framework, we show that symmetry breaking can instead enhance light trapping by strengthening in-plane nonlocal coupling pathways.
-Through diffraction and multiple-scattering analyses, we demonstrate that diffractive bands and Mie-tronic supermodes originate from the same underlying Mie resonances but differ fundamentally in physical nature.
-Finite arrays exhibit $Q$-factor enhancement driven by redistributed radiation channels, reversing the trend predicted by infinite-lattice theory.
-We further show that controlled symmetry breaking opens new electromagnetic coupling channels, enabling polarization conversion in nonlocal metasurfaces.
-These findings establish a unified wave picture linking scattering and diffraction theories and outline design principles for multifunctional metasurfaces that exploit nonlocality for advanced light manipulation, computation, and emission control.
+It is usually believed that symmetry breaking in photonic systems leads to weaker optical confinement, such as in the case of metasurfaces when bound states in the continuum are replaced by quasi-bound states with lower quality factors ($Q$ factors).
+Here we show that symmetry breaking can instead enhance light trapping by strengthening in-plane nonlocal coupling pathways.
+We consider finite-size arrays of optical resonators supporting Mie resonances (a Mie-tronics platform) and employ diffraction and multiple-scattering analyses.
+We demonstrate that diffractive bands and Mie-tronics supermodes originate from the same underlying Mie resonances but differ fundamentally in their physical nature.
+Finite arrays exhibit $Q$-factor enhancement driven by redistributed radiation channels, and reversing the trends predicted by infinite-lattice theories.
+We reveal that controlled symmetry breaking opens new electromagnetic coupling channels, enabling polarization conversion in nonlocal metasurfaces.
+These novel findings establish a unified wave-physics platform linking both scattering and diffraction theories.
+Also, they outline the design principles for multifunctional metasurfaces that exploit nonlocality for advanced light manipulation, computation, and emission control.
 preprint: APS/123-QED
 I Introduction
-Optical metasurfaces are rapidly evolving from fundamental science to practical technology, enabling compact platforms for imaging, sensing, energy harvesting, and information processing [Brongersma2025second].
+The field of optical metasurfaces is rapidly evolving from fundamental science to practical technologies, enabling compact platforms for imaging, sensing, energy harvesting, and information processing [Brongersma2025second].
 Conventional metasurfaces comprise weakly interacting subwavelength elements, allowing a local description.
@@
 He explicitly noted that Fresnel’s local approximation fails when the grating period approaches the wavelength, observing that the recesses “act as resonators” [Rayleigh1907dynamical].
-Rayleigh thus recognized, decades ahead of modern electromagnetic modeling, that resonance fundamentally alters diffraction and that a “more strictly dynamical theory” was required.
-Fano later provided this framework [Fano1941theory], introducing coupling among diffracted and evanescent orders to describe how these “Rayleigh cavities” form quasi-stationary surface waves that interfere with the continuum, giving rise to asymmetric Fano resonances.
-The Rayleigh–Fano formalism laid the foundation for modern numerical solvers such as the rigorous coupled-wave analysis (RCWA) or Fourier modal method [Liu2012s4], which we use here to calculate nonlocal diffraction bands in ideal, infinitely periodic metasurfaces.
-The design of nonlocal metasurfaces often begins with idealized infinite arrays, where symmetry breaking within the unit cell serves as a versatile tool for tailoring the dispersion and radiation properties of diffractive bands [Nguyen2018symmetry, Overvig2020selection, You2024resonance, Barreda2025bound].
+Rayleigh thus recognized, decades ahead of modern electromagnetic modeling, that resonance fundamentally alters diffraction and that a “more strictly dynamical theory” is required.
+Ugo Fano later provided this framework [Fano1941theory], introducing coupling among diffracted and evanescent orders to describe how these “Rayleigh cavities” form quasi-stationary surface waves that interfere with the continuum, giving rise to asymmetric Fano lineshapes.
+The Rayleigh–Fano formalism laid the foundation for modern numerical solvers such as the rigorous coupled-wave analysis (RCWA) or Fourier modal method [Liu2012s4], which are employed here to calculate nonlocal diffraction bands in ideal, infinitely periodic metasurfaces.
+The design of nonlocal metasurfaces often begins with idealized infinite arrays of resonators, where symmetry breaking within the unit cell serves as a versatile tool for tailoring the dispersion and radiation properties of diffractive bands [Nguyen2018symmetry, Overvig2020selection, You2024resonance, Barreda2025bound].
 Engineering such diffractive bands has recently fueled advances in analog optical computing and wave-based information processing [Guo2018photonic, Kwon2018nonlocal, Komar2021edge, Liu2024edge].
@@
 As the array size decreases, translational symmetry is broken and the momentum conservation underlying the Fourier modal method no longer holds.
-The resulting collective resonances become discrete and strongly dependent on the global geometry of the structure.
+The resulting collective resonances become discrete and strongly dependent on a global geometry of the structure.
 In this finite regime, multiple scattering and modal hybridization provide a more accurate physical description [Lagendijk1996resonant, Hoang2025collective].
-At the microscopic level, these collective effects originate from interactions among Mie modes—the localized resonances of individual dielectric or metallic elements—which constitute the building blocks of Mie-tronics [Rybin2024metaphotonics].
-Beyond governing the response of isolated particles [Li2024machine, Barati2025toward], the coupling and interference of Mie modes across arrays yield bonding and antibonding supermodes.
+At the microscopic level, the collective effects originate from interactions among Mie modes—the localized resonances of individual dielectric or metallic elements—which constitute the building blocks of Mie-tronics [Rybin2024metaphotonics].
+Beyond governing the response of isolated particles [Li2024machine, Barati2025toward], the coupling of Mie modes across arrays yield bonding and anti-bonding supermodes.
 These supermodes can enhance device performance and enable entirely new metadevice concepts [Krasnok2012all, Babicheva2024mie, Ha2024optoelectronic, Mao2025lateral].
-In this work, we systematically investigate how symmetry breaking influences both diffractive and Mie-tronic supermodes.
+In this work, we study systematically how symmetry breaking can affect both diffractive and Mie-tronics supermodes.
 Together with the accompanying Letter [Hoang2025unconventional], we show that, contrary to expectations based on infinite-array diffraction bands, symmetry breaking in finite metasurfaces can enhance in-plane multiple scattering and thereby increase the quality factor ($Q$) of certain nonlocal modes.
-We further demonstrate that coupling among Mie modes produces additional bonding and antibonding supermodes beyond the Fourier-mode couplings of classical diffraction theory, expanding the functional landscape of nonlocal metasurfaces.
-Notably, while antibonding supermodes remain robust in the presence of a quartz substrate, bonding supermodes are suppressed.
-Finally, we show that distinct unit-cell geometries can support identical Mie-resonant supermodes, underscoring the robustness and design versatility of the Mie-tronic framework for both diffractive and finite-size nonlocal metasurfaces.
+We further demonstrate that coupling among Mie modes produces additional bonding and anti-bonding supermodes beyond the Fourier-mode couplings of classical diffraction theory, expanding the functional landscape of nonlocal metasurfaces.
+Notably, while anti-bonding supermodes remain robust in the presence of a quartz substrate, bonding supermodes are suppressed.
+Finally, we show that distinct unit-cell geometries can support identical Mie-resonant supermodes, underscoring the robustness and design versatility of the Mie-tronics framework for both diffractive and finite-size nonlocal metasurfaces.
 II Fundamentals of Mie-tronics
@@
 Advances in nanoscience have since extended scattering analysis to particles of arbitrary shape [Kildishev2025art], enabling entirely new classes of photonic devices [Dorodnyy2023mie].
-This convergence of multipole physics and nanophotonics defines the emerging paradigm of Mie-tronics [Won2019into]—a framework that unifies classical scattering theory with modern metasurface design.
-II.2 Diffraction and Scattering Frameworks
+This convergence of multipole physics and nanophotonics defines the emerging paradigm of Mie-tronics [Won2019into, kivshar_Mie]—a framework that unifies classical scattering theory with modern metasurface design.
+II.2 Diffraction and scattering
 The early development of scattering theory was closely intertwined with the study of diffraction.
@@
 Within this unified framework, Mie-tronics bridges Fourier-mode analyses with multipole expansions, revealing their common physical foundations and providing a coherent language for describing nonlocal metasurfaces.
-II.3 Mie Coefficients as Building Blocks
+II.3 Mie modes as the building blocks
 Figure 1: Beyond spheres: unit cells in Mie-tronics.
@@
 Spectral overlap among Mie modes thus lays the foundation for collective optical phenomena in arrays.
-When resonant unit cells are periodically coupled, their individual multipolar responses hybridize to form delocalized collective modes—Mie-tronic supermodes—in finite arrays.
-In the following section, we examine how these collective Mie-tronic supermodes connect to Bloch bands in infinite periodic systems.
-II.4 Beyond Bloch Bands: Antibonding and Bonding Mie-tronic Supermodes
-Figure 3: Mie-tronic origin of supermodes in finite arrays and their connection to Bloch bands in photonic crystals.
-(a) Purcell-factor spectrum for the configuration in the inset, revealing two well-separated bands (antibonding and bonding) with pronounced supermodes labeled $A_{1,2}$ and $B_{1}$.
-(b),(c) Magnetic ($H_{x}$) and electric ($E_{z}$) field distributions for the antibonding mode $A_{1}$; the inset shows the corresponding far-field radiation pattern.
-(d) Bloch bands associated with the antibonding and bonding supermodes in (a).
+When resonant unit cells are periodically coupled, their individual multipolar responses hybridize to form delocalized collective modes—Mie-tronics supermodes—in finite arrays.
+In the following section, we examine how these collective Mie-tronics supermodes connect to Bloch bands in infinite periodic systems.
+II.4 Beyond Bloch bands: Anti-bonding and bonding supermodes
+Figure 3: Mie-tronics origin of supermodes in finite arrays and their connection to Bloch bands in photonic crystals.
+(a) Purcell-factor spectrum for the configuration in the inset, revealing two well-separated bands (anti-bonding and bonding) with pronounced supermodes labeled $A_{1,2}$ and $B_{1}$.
+(b),(c) Magnetic ($H_{x}$) and electric ($E_{z}$) field distributions for the anti-bonding mode $A_{1}$; the inset shows the corresponding far-field radiation pattern.
+(d) Bloch bands associated with the anti-bonding and bonding supermodes in (a).
 (e),(f) Same as (b),(c), but for the bonding mode $B_{1}$; the inset highlights dominant in-plane radiation leakage.
@@
 Two distinct bands appear, each containing well-defined supermodes labeled $A_{1}$, $A_{2}$, and $B_{1}$.
-The $A_{1}$ and $B_{1}$ modes correspond to the $\Gamma$- and $M$-point edges of the antibonding ($A$) and bonding ($B$) Bloch bands in Fig.
+The $A_{1}$ and $B_{1}$ modes correspond to the $\Gamma$- and $M$-point edges of the anti-bonding ($A$) and bonding ($B$) Bloch bands in Fig.
 3(d), which emerge in the limit of an infinite array.
@@
 As a result, the calculated band-edge wavelengths align precisely with the observed supermodes [Hoang2020collective].
-The antibonding and bonding characters of these modes are evident from their near- and far-field profiles.
-Figures 3(b,c) show the $A_{1}$ (antibonding) supermode, while Figs.
+The anti-bonding and bonding characters of these modes are evident from their near- and far-field profiles.
+Figures 3(b,c) show the $A_{1}$ (anti-bonding) supermodes, while Figs.
 3(e,f) show the $B_{1}$ (bonding) mode.
@@
 These hotspots signify stronger in-plane coupling and larger Purcell enhancement for the bonding mode.
-Such enhanced local fields enable efficient interaction with emitters or detectors oriented along the $z$ direction, suggesting on-chip architectures where Mie-tronic modes act as optical communication channels linking integrated transmitters and receivers—forming scalable photonic networks capable of emulating quantum many-body interactions.
+Such enhanced local fields enable efficient interaction with emitters or detectors oriented along the $z$ direction, suggesting on-chip architectures where Mie-tronics modes act as optical communication channels linking integrated transmitters and receivers—forming scalable photonic networks capable of emulating quantum many-body interactions.
 The topological properties of Mie modes, in this context, could underpin future classical and nonclassical simulators and light sources [Nielsen2010quantum, Berloff2017realizing, Sergeeva2025coherent, Deng2025enhancement].
@@
 In $A_{1}$, the moments are aligned, leading to repulsive interactions; in $B_{1}$, they are anti-aligned, resulting in attractive coupling.
-Put differently, bonding supermodes extend their fields into the environment, whereas antibonding modes confine them within the unit cells.
-This difference explains the greater robustness of antibonding modes to perturbations introduced by a quartz substrate, as discussed in the next section.
+Put differently, bonding supermodes extend their fields into the environment, whereas anti-bonding modes confine them within the unit cells.
+This difference explains the greater robustness of anti-bonding modes to perturbations introduced by a quartz substrate, as discussed in the next section.
 The far-field radiation patterns [insets of Figs.
 3(b,e)] provide complementary evidence.
-The bonding mode $B_{1}$ radiates mainly in-plane, while the antibonding mode $A_{1}$ emits strongly out of plane.
+The bonding mode $B_{1}$ radiates mainly in-plane, while the anti-bonding mode $A_{1}$ emits strongly out of plane.
 Similar to the hole arrays discussed in the accompanying Letter [Hoang2025unconventional], the $A_{1}$ mode produces an out-of-plane vortex beam, confirming its collective MD origin.
@@
 These symmetry considerations offer a useful guideline for tailoring diffractive bands through controlled symmetry breaking, as we demonstrate in the following section.
-Nevertheless, the collective nature of Mie-tronic supermodes allows normal-incidence excitation in sufficiently large arrays [Hoang2025collective].
+Nevertheless, the collective nature of Mie-tronics supermodes allows normal-incidence excitation in sufficiently large arrays [Hoang2025collective].
 In Fig.
@@
 The discrete peaks in Fig.
-3(a) arise from the radiative boundary conditions of finite arrays, capturing high-$Q$ supermodes inaccessible to conventional Bloch-wave analysis, which assumes in-plane periodicity and thus yields only continuous bonding and antibonding bands [Fig.
+3(a) arise from the radiative boundary conditions of finite arrays, capturing high-$Q$ supermodes inaccessible to conventional Bloch-wave analysis, which assumes in-plane periodicity and thus yields only continuous bonding and anti-bonding bands [Fig.
 3(d)].
-These finite-array effects highlight how the interplay between collective coupling and structural symmetry profoundly reshapes resonance behavior—a theme we now pursue by examining how symmetry breaking governs the formation and quality factors of Mie-tronic supermodes.
-III Impact of Symmetry Breaking on Mie-tronic Supermodes
+These finite-array effects highlight how the interplay between collective coupling and structural symmetry profoundly reshapes resonance behavior—a theme we now pursue by examining how symmetry breaking governs the formation and quality factors of Mie-tronics supermodes.
+III Impact of symmetry breaking on Mie-tronics supermodes
 Symmetry breaking in unit cells has been extensively explored for engineering diffractive bands, yet its influence on collective resonances in finite metasurfaces remains far less understood.
-Here, we show that symmetry breaking not only modifies the coupling between Mie-tronic supermodes and external radiation but can also produce counterintuitive $Q$-factor enhancement—opposite to trends predicted by infinite-lattice Bloch theory.
+Here, we show that symmetry breaking not only modifies the coupling between Mie-tronics supermodes and external radiation but can also produce counterintuitive $Q$-factor enhancement—opposite to trends predicted by infinite-lattice Bloch theory.
 Moreover, the distinct topological characteristics of Mie modes provide a natural foundation for designing nonlocal metasurfaces with tailored confinement, coupling, and polarization properties.
-III.1 Robustness of Antibonding Supermodes Against Symmetry Breaking
-Figure 4: Symmetry breaking enhances in-plane multiple scattering, preserves antibonding supermodes, and suppresses bonding counterparts in square and T-shaped arrays.
+III.1 Robustness of anti-bonding supermodes against symmetry breaking
+Figure 4: Symmetry breaking enhances in-plane multiple scattering, preserves anti-bonding supermodes, and suppresses bonding counterparts in square and T-shaped arrays.
 (a) Purcell-factor spectrum for a magnetic dipole $m_{x}$ at the center of a $9\times 9$ array of square unit cells.
-The presence of a quartz substrate ($n\simeq 1.45$) suppresses the bonding mode $B_{1}$, while the antibonding mode $A_{1}$ remains nearly unaffected.
+The presence of a quartz substrate ($n\simeq 1.45$) suppresses the bonding mode $B_{1}$, while the anti-bonding mode $A_{1}$ remains nearly unaffected.
 (b),(c) Magnetic near-field distribution and vortex-like far-field pattern of $A_{1}$.
@@
 The far-field pattern highlights enhanced in-plane multiple scattering along in-plane channels (ICs) versus suppressed out-of-plane channels (OCs).
-Figure 4 summarizes the effects of in-plane and out-of-plane symmetry breaking on Mie-tronic supermodes.
+Figure 4 summarizes the effects of in-plane and out-of-plane symmetry breaking on Mie-tronics supermodes.
 In-plane symmetry breaking is introduced by replacing square unit cells with T-shaped ones (same lattice vectors but reduced unit-cell symmetry), while out-of-plane breaking is implemented by adding a quartz substrate.
 For the free-standing square lattice, two well-separated resonant bands appear [Fig.
-4(a)], corresponding to the antibonding ($A_{1}$) and bonding ($B_{1}$) supermodes at the band edge.
+4(a)], corresponding to the anti-bonding ($A_{1}$) and bonding ($B_{1}$) supermodes at the band edge.
 The magnetic-field distribution of $A_{1}$ [Fig.
@@
 Introducing a substrate strongly perturbs the bonding mode $B_{1}$, whose attractive field distribution extends into the environment and overlaps with substrate-supported channels.
-In contrast, the antibonding mode remains largely unaffected owing to its repulsive character, which minimizes near-field interaction with the substrate.
+In contrast, the anti-bonding mode remains largely unaffected owing to its repulsive character, which minimizes near-field interaction with the substrate.
 The vortex-like far-field pattern of $A_{1}$ [Fig.
@@
 5.
-III.2 $Q$ Enhancement via Symmetry Breaking
+III.2 $Q$-enhancement via symmetry breaking
 Figure 5: Symmetry breaking lowers the $Q$ of Bloch modes in infinite lattices but enhances the $Q$ of supermodes in finite metasurface arrays.
-(a) Antibonding Bloch bands above the light line for the four unit-cell types.
+(a) anti-bonding Bloch bands above the light line for the four unit-cell types.
 (b) $Q$ of Bloch modes at $\Gamma$: symmetry breaking reduces the divergent $Q$ of the square unit to a finite value in the T-shaped unit, while sphere and hole lattices show much higher $Q$.
 (c) $Q$ of supermodes in finite arrays: symmetry breaking boosts the $Q$ of T-shaped arrays beyond that of square arrays for sizes $5\times 5$ to $33\times 33$, with both outperforming sphere and hole arrays.
-Figure 5(a) shows the antibonding Bloch bands above the light line for the four unit-cell geometries.
+Figure 5(a) shows the anti-bonding Bloch bands above the light line for the four unit-cell geometries.
 The calculated band-edge wavelengths agree well with the corresponding $A_{1}$ supermodes (the hole-array case is detailed in the accompanying Letter [Hoang2025unconventional]).
@@
 Although optical analogues of Wannier functions have been proposed for modeling light localization, the exponential confinement intrinsic to electronic Wannier functions is not directly applicable to electromagnetic waves.
-As shown in the accompanying Letter [Hoang2025unconventional], multipole functions provide a more natural and physically rigorous basis for investigating light localization within the Mie-tronic framework.
+As shown in the accompanying Letter [Hoang2025unconventional], multipole functions provide a more natural and physically rigorous basis for investigating light localization within the Mie-tronics framework.
 IV Diffractive Nonlocal Metasurfaces
@@
 Their ability to sustain collective, system-scale optical modes enables both deep miniaturization of on-chip architectures and performance that surpasses conventional optical components [Guo2020squeeze, Miller2023optics, Wan2024multidimensional, Javidi2025roadmap].
-IV.1 Mie-Resonant Origin of Nonlocal Diffractive Bands
+IV.1 Mie-resonance origin of nonlocal diffractive bands
 Figure 6: Mie-resonant origins of diffractive nonlocal bands.
@@
 Consequently, most analyses—including the present one—focus on the two lowest-order multipoles, MD and ED, whose well-separated resonances allow controlled dispersion engineering through unit-cell geometry.
-In coupled arrays, both dipole and quadrupole modes split into bonding and antibonding branches.
-Within diffraction theory, only the antibonding modes couple efficiently to the radiation continua, giving rise to observable diffractive bands.
+In coupled arrays, both dipole and quadrupole modes split into bonding and anti-bonding branches.
+Within diffraction theory, only the anti-bonding modes couple efficiently to the radiation continua, giving rise to observable diffractive bands.
 Because dipole and quadrupole resonances occur at relatively low frequencies, their associated diffraction bands are spectrally well separated and tunable through geometry.
@@
 This symmetry argument explains the polarization selectivity of the MD resonance and clarifies the physical origin of symmetry-protected BICs in diffractive nonlocal metasurfaces.
-More fundamentally, the symmetry mismatch arises from destructive interference of multipolar displacement currents within the symmetric unit cell, which cancels the radiation of the MD supermode along the surface normal [Hoang2025collective].
-This interference picture also explains how symmetry breaking restores coupling between P-polarized plane waves and the MD supermode, as demonstrated in the following subsection.
-IV.2 Symmetry Breaking Enables Polarization Conversion
+More fundamentally, the symmetry mismatch arises from destructive interference of multipolar displacement currents within the symmetric unit cell, which cancels the radiation of the MD supermodes along the surface normal [Hoang2025collective].
+This interference picture also explains how symmetry breaking restores coupling between P-polarized plane waves and the MD supermodes, as demonstrated in the following subsection.
+IV.2 Polarization conversion via symmetry breaking
 Figure 7: Symmetry breaking enables polarization conversion in nonlocal metasurfaces.
@@
 These results show that symmetry considerations within the Mie-tronics framework provide a robust strategy for designing nonlocal metasurfaces with tunable polarization filtering and conversion.
-V Discussion and Conclusion
-Metasurfaces have become a cornerstone of modern photonics, offering miniaturized, chip-scale platforms for advanced optical functionalities.
+V Discussion and Conclusions
+Metasurfaces have become a cornerstone of modern photonics, offering future miniaturized, chip-scale platforms for advanced optical functionalities.
 Yet, pushing their performance beyond current limits requires explicit consideration of nonlocal interactions among unit cells.
-Within the Mie-tronics framework, we have shown that symmetry breaking enables precise control over the coupling between metasurface supermodes and external excitations, leading to polarization conversion and enhanced light confinement in finite metastructures.
+Within the Mie-tronics framework, we have shown that symmetry breaking enables precise control over the coupling between metasurface supermodes and external excitations, leading to polarization conversion and enhanced light confinement in finite-size metastructures.
 These mechanisms lay the foundation for emerging functionalities such as analog optical computing and on-chip information processing.
@@
 This research is supported by the National Research Foundation, Singapore, and A*STAR under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0009), the Australian Research Council (Grant No.
-DP210101292), and International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (Contract FA520923C0023).
+DP210101292).
 D.L.